OpenStudy (anonymous):

45mins/2mi=?h/1mi again I can't do this its equivalent rate

OpenStudy (anonymous):

are you looking for a speed? Mph?

OpenStudy (anonymous):

45/2=x/1 x=22.5 minutes 22.5 minutes divided by 60 gives you .375 hours for every 1 mile traveled.

OpenStudy (anonymous):

Well it shows ?h/1mi